suppose you invest $5,000 into a project and after 17 months yiu sell your share in this project for $7,000
A) Calculate the interest rate earned per month on this investment.

B) Calculate the effective annual yield of this investment
Answer #1

a)

rate of interest (monthly) = r

t = 17 months , FV = $ 7000 and PV = $ 5000

5000 x (1 + r)17 = 7000

r = 2%

b)

effective annual yield = (1 + r)12 - 1 = 26.81%